# Technical Report

### v1

@@
Technical Report
XMainframe: A Large Language Model for
Mainframe Modernization
Anh T. V. Dau, Hieu Trung Dao, Anh Tuan Nguyen, Hieu Trung Tran
Phong X. Nguyen, Nghi D. Q. Bui
FPT Software AI Center
Vietnam
https://github.com/FSoft-AI4Code/XMainframe
@@ 1 Introduction
Large Language Models for code (CodeLLMs) excel in processing and understanding source code across various programming languages such as Python, C++, Java, C#, Rust, Go, etc., as well as descriptive texts Qin et al. (2023); Touvron et al. (2023); Roziere et al. (2023); Jiang et al. (2024); Team (2024); Manh et al. (2023); Zheng et al. (2024); Li et al. (2023); Wang et al. (2023b); Feng et al. (2020); Wang et al. (2021); Bui et al. (2023).. Their ability to recognize patterns, syntax, and semantics makes them highly effective at tasks such as code completion, bug detection, and generating human-readable explanations. These models can bridge the gap between code and documentation by comprehending and generating natural language descriptions
Mainframe Modernization: Mainframe software systems are crucial to the daily operations of many of the world’s largest corporations, including numerous Fortune 1000 companies IBM . These systems are used extensively in domains such as banking, finance, and government, where they manage large-scale user bases and applications. Despite their origins in the 1950s, COBOL (Common Business Oriented Language) remains widely used in mainframe applications. It is estimated that over 220 billion lines of COBOL code are currently in use, with 1.5 billion lines written annually Taulli (2020). Additionally, COBOL systems manage USD 3 trillion in commerce daily Cassel (2017). However, the retirement of many COBOL developers and mainframe experts poses a significant challenge for maintaining and modernizing these systems. In 2014, American Banker reported that banks face difficulties in attracting young tech talent and there is a shortage of professionals with mainframe and COBOL skills Crosman (2014). This highlights the urgent need for innovative solutions to bridge the gap between legacy COBOL systems and modern technologies, denoted as mainframe modernization. There is recent interest in adapting mainstream CodeLLMs to modernize legacy systems written in aging languages like COBOL into modern languages such as C++ and Java to address the shortage and retirement of COBOL developers and mainframe experts.
@@ 2.1 Code Large Language Models
There have been numerous Code-LLMs trained on massive datasets to achieve remarkable advancements across diverse coding tasks, including code generation Roziere et al. (2023); Touvron et al. (2023); Li et al. (2023); Jiang et al. (2024); Feng et al. (2020), code summarization Ahmed and Devanbu (2022); Lu et al. (2021); Gao et al. (2023); Su and McMillan (2024), program repair Xia and Zhang (2022); Wei et al. (2023); Xia et al. (2023), etc. Besides, LLMs showcase unforeseen capabilities that allow them to adapt to different domains without modifying their parameters by using discrete prompting. These prompts created by humans or by LLMs themselves involve providing instructions along with relevant contexts as input to refine the generation Luo et al. (2023); Wang et al. (2023a). Related to instruction tuning is chain-of-thought promoting, where models are encouraged to explain the reasoning when given a complex problem, thereby increasing the likelihood that their final answer is correct Wei et al. (2022). Recently, several studies proposed multi-agent collaborations, where each agent focuses on a unique task, such as code generation or task planning, to enhance the effectiveness of LLM-based code generation Chen et al. (2023); Qian et al. (2023); Huang et al. (2023).
@@ 2.2 LLMs for domain-specific tasks
Code-related datasets have been introduced to verify empirical research across various programming languages and tackle challenges in multiple areas of software engineering Odena et al. (2021); Iyer et al. (2018); Chen et al. (2021); Nguyen et al. (2023). However, low-resource language such as COBOL has not received enough attention from scientific and academic researchers. Therefore, a big barrier exists to training LLMs for COBOL on a large scale. OpenCBS Lee et al. (2022) is one of the pioneering efforts leveraging public forums to create a COBOL set for defect detection. Another dataset, X-COBOL Ali et al. (2023), comprises 84 COBOL repositories collected from GitHub. Despite undergoing a data extraction pipeline, this dataset falls short of the intended quality as the authors relied on GitHub stars for filtering the repositories, which is not a sufficient metric for determining the quality of a repository Borges and Valente (2018). More recently, BloopAI announced the COBOLEval Gordon-Hall (2024), a benchmark designed to evaluate legacy code completion tasks. It consists of 146 coding problems converted into COBOL from HumanEval benchmark Chen et al. (2021), which is a Python dataset. Nevertheless, it is unrealistic since COBOL is primarily used for business and finance systems, not for solving programming challenges. To the best of our knowledge, there is no dataset that covers diverse tasks related to COBOL language and legacy systems.
@@ 3.2 Dataset for Model Instruct
Besides using LLMs to solve tasks, recent works have treated LLMs as data generators Ye et al. (2022, 2023); Yu et al. (2024). With only a few examples, LLMs are able to generate more high-quality data through in-context learning and prompting. The experiments showed that task-specific models trained on generated data can beat the performance of original LLMs while maintaining a low inference cost. Inspired by the self-instruct approach Ouyang et al. (2022), we further enrich Mainframe-Instruct from the seed data by harnessing the capabilities of popular LLMs, including OpenAI GPT-3.5-turbo, Mistral-Instruct 7B Jiang et al. (2023), Neural-Chat 7B Intel (2023), and Mixtral-Instruct 8x7B model Jiang et al. (2024), which are trained on numerous languages and achieved high performance on various NLP and software benchmarks. Below are the examples of prompts that we use to generate data from a sub-topic and seed data:
@@ 4.3 Model Upscale
Inspired by Kim et al. (2023), we employ the depth up-scaling method to expand the base model without introducing additional modules or dynamic expert selection methods like Mixture of Experts (MoE) Shazeer et al. (2017); Komatsuzaki et al. (2022). This approach maintains high efficiency during both training and inference. The depthwise scaling process, illustrated in Figure 6, involves two steps: expanding the base model and continuing pretraining. First, the base model, consisting of $n$ layers, is duplicated. We then remove the last $m$ layers from the original model and the first $m$ layers from its duplicate, creating two separate models with nmn-m layers each. These parts are combined to form a scaled model with s=2(nm)s=2(n-m) layers. For our purposes, we choose $m=6$. With $n=30,m=6,s=48$, this process is depicted in Step 1 of Figure 6. As a result, we scale up DeepSeek-Coder 7B, originally with 30 layers, to a 10.5B model with 48 layers.

### v2

# XMainframe: A Large Language Model for Mainframe Modernization

@@
Abstract
@@ 1 Introduction
Large Language Models for code (CodeLLMs) excel in processing and understanding source code across various programming languages such as Python, C++, Java, C#, Rust, Go, etc., as well as descriptive texts Qin et al. (2023); Touvron et al. (2023); Roziere et al. (2023); Jiang et al. (2024); Team (2024); Manh et al. (2023); Zheng et al. (2024); Li et al. (2023); Wang et al. (2023b); Feng et al. (2020); Wang et al. (2021); Bui et al. (2023). Their ability to recognize patterns, syntax, and semantics makes them highly effective at tasks such as code completion, bug detection, and generating human-readable explanations. These models can bridge the gap between code and documentation by comprehending and generating natural language descriptions
Mainframe Modernization: Mainframe software systems are crucial to the daily operations of many of the world’s largest corporations, including numerous Fortune 1000 companies. These systems are used extensively in domains such as banking, finance, and government, where they manage large-scale user bases and applications. Despite their origins in the 1950s, COBOL (Common Business Oriented Language) remains widely used in mainframe applications. It is estimated that over 220 billion lines of COBOL code are currently in use, with 1.5 billion lines written annually Taulli (2020). Additionally, COBOL systems manage USD 3 trillion in commerce daily Cassel (2017). However, the retirement of many COBOL developers and mainframe experts poses a significant challenge for maintaining and modernizing these systems. In 2014, American Banker reported that banks face difficulties in attracting young tech talent and there is a shortage of professionals with mainframe and COBOL skills Crosman (2014). This highlights the urgent need for innovative solutions to bridge the gap between legacy COBOL systems and modern technologies, denoted as mainframe modernization. There is recent interest in adapting mainstream CodeLLMs to modernize legacy systems written in aging languages like COBOL into modern languages such as C++ and Java to address the shortage and retirement of COBOL developers and mainframe experts.
@@ 2.1 Code Large Language Models
There have been numerous Code-LLMs trained on massive datasets to achieve remarkable advancements across diverse coding tasks, including code generation Roziere et al. (2023); Touvron et al. (2023); Li et al. (2023); Jiang et al. (2024); Feng et al. (2020), code summarization Ahmed & Devanbu (2022); Lu et al. (2021); Gao et al. (2023); Su & McMillan (2024), program repair Xia & Zhang (2022); Wei et al. (2023); Xia et al. (2023), etc. Besides, LLMs showcase unforeseen capabilities that allow them to adapt to different domains without modifying their parameters by using discrete prompting. These prompts created by humans or by LLMs themselves involve providing instructions along with relevant contexts as input to refine the generation Luo et al. (2023); Wang et al. (2023a). Related to instruction tuning is chain-of-thought promoting, where models are encouraged to explain the reasoning when given a complex problem, thereby increasing the likelihood that their final answer is correct Wei et al. (2022). Recently, several studies proposed multi-agent collaborations, where each agent focuses on a unique task, such as code generation or task planning, to enhance the effectiveness of LLM-based code generation Chen et al. (2023); Qian et al. (2023); Huang et al. (2023).
@@ 2.2 LLMs for domain-specific tasks
Code-related datasets have been introduced to verify empirical research across various programming languages and tackle challenges in multiple areas of software engineering Odena et al. (2021); Iyer et al. (2018); Chen et al. (2021); Nguyen et al. (2023). However, low-resource language such as COBOL has not received enough attention from scientific and academic researchers. Therefore, a big barrier exists to training LLMs for COBOL on a large scale. OpenCBS Lee et al. (2022) is one of the pioneering efforts leveraging public forums to create a COBOL set for defect detection. Another dataset, X-COBOL Ali et al. (2023), comprises 84 COBOL repositories collected from GitHub. Despite undergoing a data extraction pipeline, this dataset falls short of the intended quality as the authors relied on GitHub stars for filtering the repositories, which is not a sufficient metric for determining the quality of a repository Borges & Valente (2018). More recently, BloopAI announced the COBOLEval Gordon-Hall (2024), a benchmark designed to evaluate legacy code completion tasks. It consists of 146 coding problems converted into COBOL from HumanEval benchmark Chen et al. (2021), which is a Python dataset. Nevertheless, it is unrealistic since COBOL is primarily used for business and finance systems, not for solving programming challenges. To the best of our knowledge, there is no dataset that covers diverse tasks related to COBOL language and legacy systems.
@@ 3.2 Dataset for Model Instruct
Besides using LLMs to solve tasks, recent works have treated LLMs as data generators Ye et al. (2022; 2023); Yu et al. (2024). With only a few examples, LLMs are able to generate more high-quality data through in-context learning and prompting. The experiments showed that task-specific models trained on generated data can beat the performance of original LLMs while maintaining a low inference cost. Inspired by the self-instruct approach Ouyang et al. (2022), we further enrich Mainframe-Instruct from the seed data by harnessing the capabilities of popular LLMs, including OpenAI GPT-3.5-turbo, Mistral-Instruct 7B Jiang et al. (2023), Neural-Chat 7B Intel (2023), and Mixtral-Instruct 8x7B model Jiang et al. (2024), which are trained on numerous languages and achieved high performance on various NLP and software benchmarks. Below are the examples of prompts that we use to generate data from a sub-topic and seed data:
@@ 4.3 Model Upscale
Inspired by Kim et al. (2023), we employ the depth up-scaling method to expand the base model without introducing additional modules or dynamic expert selection methods like Mixture of Experts (MoE) Shazeer et al. (2017); Komatsuzaki et al. (2022). This approach maintains high efficiency during both training and inference. The depthwise scaling process, illustrated in Figure 6, involves two steps: expanding the base model and continuing pretraining. First, the base model, consisting of $n$ layers, is duplicated. We then remove the last $m$ layers from the original model and the first $m$ layers from its duplicate, creating two separate models with $n-m$ layers each. These parts are combined to form a scaled model with $s=2(n-m)$ layers. For our purposes, we choose $m=6$. With $n=30,m=6,s=48$, this process is depicted in Step 1 of Figure 6. As a result, we scale up DeepSeek-Coder 7B, originally with 30 layers, to a 10.5B model with 48 layers.